# Modeling the Multivariate Relationship with Contextualized Representations for Effective Human-Object Interaction Detection

### v1

@@ Introduction
CATN (CVPR 2022) R50 31.86 25.15 33.84 34.44 27.69 36.45 60.1 -
GEN-VLKT (CVPR 2022) R50+ViT-B/32 33.75 29.25 35.10 36.78 32.75 37.99 62.4 64.5
ERNet (TIP 2023) EfficientNetV2-L 34.25 28.70 36.33 - - - 61.6 -
SG2HOI+ (TIP 2023) R50+ViT-B/32 33.14 29.27 35.72 35.73 32.01 36.43 63.6 65.2
Multi-Step (ACM MM 2023) R101 34.42 30.03 35.73 37.71 33.74 38.89 61.3 67.0
HOICLIP (CVPR 2023) R50+ViT-B/32 34.69 31.12 35.74 37.61 34.47 38.54 63.5 64.8
CEFA (ACM MM 2024) R50+ViT-B/32 35.00 32.30 35.81 38.23 35.62 39.02 63.5 -
DP-ADN (AAAI 2024) R50+ViT-B/32 35.91 35.82 35.44 38.99 39.61 38.80 62.6 64.8
@@
ADA-CM (ICCV 2023) R50+ViT-L/14 38.40 37.52 38.66 - - - 58.6 64.0
CLIP4HOI (NeurIPS 2023) R50+ViT-B/16 35.33 33.95 35.74 - - - - 66.3
CMMP (ECCV 2024) R50+ViT-L/14 38.14 37.75 38.25 - - - - 64.0
Pose-Aware (CVPR 2024) R50 35.86 32.48 36.86 39.48 36.10 40.49 61.1 66.6
EZ-HOI (NeurIPS 2024) R50+ViT-L/14 38.61 37.70 38.90 - - - 60.5 66.2
LAIN (CVPR 2025) R50+ViT-B/16 36.02 35.70 36.11 - - - - 65.1
HOLa (ICCV 2025) R50+ViT-L/14 39.05 38.66 39.17 - - - 60.3 66.0
CRLN-B (Ours) R50+ViT-B/16 36.70 35.16 37.16 40.17 39.02 40.51 60.2 65.9
CRLN-L (Ours) R50+ViT-L/14 39.99 40.67 39.78 43.35 44.43 43.02 60.9 66.9
The released code of PViC for V-COCO is no longer available, thus the results reproduced in (Wu et al. 2024) are reported instead.
Table 1: Comparison with state-of-the-art methods on HICO-Det and V-COCO. Bold and underline items indicate the best and second-best results, respectively. One-stage and two-stage methods are highlighted separately.
Experiments
Experiment Setup
Dataset. We evaluate our model on two widely used benchmarks, HICO-DET (Chao et al. 2018) and V-COCO (Gupta and Malik 2015). HICO-DET consists of 37,633 training and 9,546 test images, covering 600 HOI categories derived from 80 object and 117 action classes, split into 138 rare and 462 non-rare categories. V-COCO is derived from COCO (Lin et al. 2014) and contains 10,326 images (5,400 for training, 4,964 for testing) with annotations for 80 object and 24 action categories.
Implementation Details. The pre-trained DETR model with a ResNet50 (He et al. 2016) backbone is selected as our object detector. AdamW (Loshchilov 2017) is used as the optimizer, with both the learning rate and weight decay being $10^{-4}$. We set $\lambda$ to 1 during training and 2.8 during inference. Unless otherwise specified, all models are trained for 15 epochs, with a learning rate drop by a factor of 5 at the $10^{th}$ epoch. The visual encoder is based on ViT-B/16 and ViT-L/14 CLIP, and during training, the parameters of CLIP remain frozen. In the MRM module, we use two blocks for both the binary and ternary decoders and set the weighting coefficient $\alpha$ to 1. [[PREFIX]] token is manually designed, with [[ACT]] token length set to 4. All experiments are conducted on 8 NVIDIA 4090 GPUs and the batch size is 16.
Comparisons to the State-of-The-Art Methods
The experimental results on the HICO-Det and V-COCO datasets are presented in Table 1. We compared with both one-stage methods, such as CATN (Dong et al. 2022), ERNet (Lim et al. 2023), SGHOI+ (He et al. 2023), etc., and two-stage methods, like CMMP (Lei et al. 2024), EZ-HOI (Lei, Wang, and Tan 2024), LAIN (Kim, Jung, and Cho 2025), HOLa (Lei, Wang, and Tan 2025).
For the HICO-DET dataset, our proposed model demonstrates remarkable performance, outperforming the most recent work HOLa by a margin of 0.94 mAP for full categories. Notably, the performance improvement is most significant in the rare categories, where our model outperforms CMMP and EZ-HOI by margins of 2.92 mAP and 2.97 mAP, respectively. For the V-COCO dataset, our model achieves 60.9 and 66.9 role AP in S1 and S2, surpassing recent work HOLa by margins of 0.7 mAP and 0.9 mAP, respectively.
To further assess the impact of tool affordance-guided interaction modeling, we construct a benchmark subset from HICO-Det, named HICO-Det-HTO. This subset exclusively includes interactions where explicit tool usage is involved. The final subset includes 783 test images. A category-wise accuracy analysis for each action is conducted on this new dataset, allowing a comparison with the baseline method PViC (Zhang et al. 2023). As shown in Fig. 3, it can be observed that our model outperforms PViC on several actions, including “cook”, “cut”, “dry”, “fill”, and “hit”.
Figure 3: Comparison of per-category accuracy between CRLN-B and PViC (Zhang et al. 2023) on HICO-Det-HTO.
@@
Table 2: Ablation study of our model. UA: unary association, TR: ternary relationship, GB: global block, RB: regional block. $\checkmark$ indicates that the module is used.
@@
Table 3: Ablative experiments for hyper-parameters in CPL. Manual: hand-crafted prompt.
Ablations Study
@@
To demonstrate the effectiveness of our framework, we conducted several ablation studies on the HICO-Det dataset. Noting that our baseline is built on the inferior Variant E3 of PViC (Zhang et al. 2023), which uses a ResNet-50 C5 backbone without an additional feature head. By replacing the decoder’s original input feature with CLIP visual features $\mathbf{V}_{e}$, this enhanced baseline achieves higher mAP scores compared to PViC, i.e. 35.45, 33.66 and 35.99 in full, rare and non-rare settings respectively.
@@
Network Architecture Design. As shown in Table 2, the introduction of Multivariate Relationship Modeling (MRM) consistently delivers performance gains. Concretely, it leads to 0.55 (full), 0.41 (rare) and 0.59 (non-rare) mAP improvements. In contrast, Contextualized Prompt Learning (CPL) achieves promising mAP increases in the full (0.94) and non-rare (1.23) settings. However, there is a slight decrease of 0.07 in mAP for the rare category. Fortunately, it can be effectively compensated by integrating the MRM module. When combined, they form a comprehensive model that delivers outstanding overall performance, notably achieving an impressive mAP boost of 1.50 in rare cases.
@@
Multivariate Relationship Modeling. The modeled binary connection is inherently included in the baseline, which cannot be removed. Therefore, we focus on unary association (UA) and ternary relationship (TR) modeling. As shown in the “MRM” column in Table 2, both are effective in improving interaction recognition. For the Full, Rare and Non-rare metrics, associating the textual information in unary modeling can improve the mAP by 0.31, 0.44 and 0.26, respectively. Meanwhile, the sole use of ternary relationship modeling boost the mAP by 0.49, 0.34, and 0.53. The best performance is achieved when both of them are combined.
@@
Contextualized Prompt Learning. The proposed Contextual Decoder is crucial in CPL to refine textual features derived from learnable prompts. It comprises two blocks, namely the global block (GB) and the regional block (RB). As illustrated in the “CPL” column in Table 2, both layers individually enhance HOI performance, and their combined use results in the most significant improvements.
@@
Moreover, those predefined parameters in CPL also need to be carefully analyzed. In Table 3 (a), we vary the length of learnable tokens [[ACT]] in our prompt template “A photo of a person [[ACT]] {object}”. The best performance is achieved with 4 tokens, while using more than this could potentially increase the complexity of learning. Instead of using the hand-crafted [[PREFIX]] tokens like “a photo of a”, we can use learnable ones, resulting in the format “[[PREFIX]] person [[ACT]] {object}”. As shown in Table 3 (b), switching to learnable [[PREFIX]] tokens appears to be less effective. This suggests that having too much flexibility in prompts might be suboptimal, aligning with the findings presented in Table 3 (a). Finally, as given in Table 3 (c), the performance is not particularly sensitive to changes in the CPL weight (i.e. $\beta$).
@@
Figure 4: Qualitative results on HICO-Det test set. Bounding boxes of humans and objects are drawn in blue and green colors. N/I denotes no interaction.
Qualitative Results
The qualitative comparisons between our model (CRLN-B) and PViC on the HICO-Det dataset are presented in Fig. 4. Our model demonstrates a greater capability to predict actions as affordance-guided interactions (e.g. hit and fill) in scenarios where tools are essential for performing these actions. To delve deeper into this aspect, we also visualize corresponding cross-attention maps in Fig. 5. Our method tends to emphasize tool categories like bats and bottles, marked in red boxes. Notably, even though the pretrained object detector does not recognize brushes, since “paintbrush” is absent from the 80 COCO categories, our model successfully highlighted it as shown in Fig. 5. Additionally, as shown in Fig. 4, our method gives a higher confidence score for the correct category “paint” compared to PViC, which sparks our model’s potential in open-world scenarios.
@@
Figure 5: The visualization of cross-attention maps on images from Fig. 4. The tools are marked in red boxes.
Conclusion
In this paper, we have proposed CRLN, a two-stage HOI detection framework that addresses two key limitations of existing methods: the lack of tool-mediated modeling and the inability to integrate visual features into prompt learning, limiting their capacity to capture context-dependent relationships. Our approach has shown notable effectiveness in challenging HOI scenarios with complex tool-assisted interactions and ambiguous contextual cues, highlighting the benefits of multivariate relationship modeling and contextualized prompt learning. In future work, we plan to explore fine-grained affordance-guided modeling and broaden generalization to open-world HOI scenarios.
Appendix A Binary Positional Embedding
We follow the pairwise spatial encoding strategy adopted in prior work (Zhang, Campbell, and Gould 2022) to construct the binary positional embedding. Given a bounding box pair $(\mathbf{b}_{i},\mathbf{b}_{j})$, we compute a set of geometric attributes that capture their relative spatial configuration. These attributes include the normalized center coordinates, width, height, area, aspect ratio of both boxes, their intersection-over-union (IoU), as well as relative area and directional distance indicators. The computed features are concatenated together with their logarithmic transformations, resulting in a 36-dimensional vector representation. This spatial encoding is subsequently passed through a multi-layer perceptron (MLP) to obtain the final binary positional embedding used in our framework.
Appendix B Decoder Architecture
As shown in Fig. 6, we describe the structural details of DETR’s decoder as illustrated in the bottom-right part of Fig. 2 in the main paper.
Figure 6: The structure of the DETR’s decoder, which comprises three key components: a self-attention module, a cross-attention module, and a feed-forward network (FFN).
Appendix C Model Complexity
@@
Considering that the introduction of the ternary decoder and contextual decoder introduces additional trainable parameters, we conduct a comprehensive comparison with several recent one-stage and two-stage HOI models. Specifically, we compare our method with EZ-HOI (Lei, Wang, and Tan 2024), CMMP (Lei et al. 2024), and ADA-CM (Lei et al. 2023), all of which adopt ViT-L/14 as the backbone. We also include comparisons with our baseline PViC (Zhang et al. 2023) and the one-stage method HOICLIP (Ning et al. 2023).

### v2

@@ III-D Training and Inference
TABLE I: Comparison with state-of-the-art methods on HICO-Det and V-COCO. Bold and underline items indicate the best and second-best results, respectively. One-stage and two-stage methods are highlighted separately.
@@
CATN[39] (ICCV’21) R50 31.86 25.15 33.84 34.44 27.69 36.45 60.1 -
GEN-VLKT [32] (CVPR’2022) R50+ViT-B/32 33.75 29.25 35.10 36.78 32.75 37.99 62.4 64.5
ERNet [40] (TIP’2023) EfficientNetV2-L 34.25 28.70 36.33 - - - 61.6 -
SG2HOI+ [41] (TIP’2023) R50+ViT-B/32 33.14 29.27 35.72 35.73 32.01 36.43 63.6 65.2
HODN [20] (TMM’2023) R50 33.14 28.54 34.52 35.86 31.18 37.26 67.0 69.1
Multi-Step [42] (ACM MM’2023) R101 34.42 30.03 35.73 37.71 33.74 38.89 61.3 67.0
HOICLIP [34] (CVPR’2023) R50+ViT-B/32 34.69 31.12 35.74 37.61 34.47 38.54 63.5 64.8
CEFA [23] (ACM MM’2024) R50+ViT-B/32 35.00 32.30 35.81 38.23 35.62 39.02 63.5 -
DP-ADN [24] (AAAI’2024) R50+ViT-B/32 35.91 35.82 35.44 38.99 39.61 38.80 62.6 64.8
@@
ADA-CM [35] (ICCV’2023) R50+ViT-L/14 38.40 37.52 38.66 - - - 58.6 64.0
CLIP4HOI [36] (NeurIPS’2023) R50+ViT-B/16 35.33 33.95 35.74 - - - - 66.3
CMMP [14] (ECCV’2024) R50+ViT-L/14 38.14 37.75 38.25 - - - - 64.0
Pose-Aware [7] (CVPR’2024) R50 35.86 32.48 36.86 39.48 36.10 40.49 61.1 66.6
EZ-HOI [15] (NeurIPS’2024) R50+ViT-L/14 38.61 37.70 38.90 - - - 60.5 66.2
LAIN [43] (CVPR’2025) R50+ViT-B/16 36.02 35.70 36.11 - - - - 65.1
HOLa [44] (ICCV’2025) R50+ViT-L/14 39.05 38.66 39.17 - - - 60.3 66.0
CRL-B (Ours) R50+ViT-B/16 36.70 35.16 37.16 40.17 39.02 40.51 60.2 65.9
CRL-L (Ours) R50+ViT-L/14 39.99 40.67 39.78 43.35 44.43 43.02 60.9 66.9
The released code of PViC for V-COCO is no longer available, thus the results reproduced in [7] are reported instead.
TABLE II: Ablation study of our model. UA: unary association, TR: ternary relationship, GB: global block, RB: regional block. $\checkmark$ indicates that the module is used.
@@
IV Experiments
IV-A Experiment Setup
Dataset. We evaluate our model on two widely used benchmarks, HICO-DET [45] and V-COCO [46]. HICO-DET consists of 37,633 training and 9,546 test images, covering 600 HOI categories derived from 80 object and 117 action classes, split into 138 rare and 462 non-rare categories. V-COCO is derived from COCO [47] and contains 10,326 images (5,400 for training, 4,964 for testing) with annotations for 80 object and 24 action categories.
Zero-shot Setting. Following prior works [34, 15], we conduct our zero-shot experiments under four distinct configurations: Rare First Unseen Combination (RF-UC), Non-rare First Unseen Combination (NF-UC), Unseen Verb (UV), and Unseen Object (UO). In the RF-UC setting, we select tail HOI categories as unseen categories, while in the NF-UC setting, we use head HOI categories as unseen categories. Under the UV and UO settings, some verb or object categories are not included in the training set, respectively.
Implementation Details. The pre-trained DETR model with a ResNet50 [48] backbone is selected as our object detector. AdamW [49] is used as the optimizer, with both the learning rate and weight decay being $10^{-4}$. We set $\lambda$ to 1 during training and 2.8 during inference. Unless otherwise specified, all models are trained for 15 epochs, with a learning rate drop by a factor of 5 at the $10^{th}$ epoch. The visual encoder is based on ViT-B/16 and ViT-L/14 CLIP, and during training, the parameters of CLIP remain frozen. In the MRM module, we use two blocks for both the binary and ternary decoders and set the weighting coefficient $\alpha$ to 1. [[PREFIX]] token is manually designed, with [[ACT]] token length set to 4. All experiments are conducted on 8 NVIDIA 4090 GPUs and the batch size is 16. The computational environment runs Ubuntu 22.04, with Python version 3.7, PyTorch version 1.10.0, torchvision version 0.11.0, and CUDA version 11.3.
IV-B Comparison to the State-of-The-Art
The experimental results on the HICO-Det and V-COCO datasets are presented in Table I. We compared with both one-stage methods, such as CATN [39], ERNet [40], SGHOI+ [41], etc., and two-stage methods, like CMMP [14], EZ-HOI [15], LAIN [43], HOLa [44]. For the HICO-DET dataset, our proposed model demonstrates remarkable performance, outperforming the most recent work HOLa by a margin of 0.94 mAP for full categories. Notably, the performance improvement is most significant in the rare categories, where our model outperforms CMMP and EZ-HOI by margins of 2.92 mAP and 2.97 mAP, respectively. For the V-COCO dataset, our model achieves 60.9 and 66.9 role AP in S1 and S2, surpassing recent work HOLa by margins of 0.7 mAP and 0.9 mAP, respectively.
To further assess the impact of tool affordance-guided interaction modeling, we construct a benchmark subset from HICO-Det, named HICO-Det-HTO. This subset exclusively includes interactions where explicit tool usage is involved. The final subset includes 783 test images. A category-wise accuracy analysis for each action is conducted on this new dataset, allowing a comparison with the baseline method PViC [31]. As shown in Fig. 3, it can be observed that our model outperforms PViC on several actions, including “cook”, “cut”, “dry”, “fill”, and “hit”.
Figure 3: Comparison of per-category accuracy between CRL-B and PViC [31] on HICO-Det-HTO.
TABLE III: Ablative experiments for hyper-parameters in CPL. Manual: hand-crafted prompt.
@@ IV-B Comparison to the State-of-The-Art
IV-C Ablations Study
@@
To demonstrate the effectiveness of our framework, we conducted several ablation studies on the HICO-Det dataset. Noting that our baseline is built on the inferior Variant E3 of PViC [31], which uses a ResNet-50 C5 backbone without an additional feature head. By replacing the decoder’s original input feature with CLIP visual features $\mathbf{V}_{e}$, this enhanced baseline achieves higher mAP scores compared to PViC, i.e., 35.45, 33.66 and 35.99 in full, rare and non-rare settings respectively.
@@
Network Architecture Design. As shown in Table II, the introduction of Multivariate Relationship Modeling (MRM) consistently delivers performance gains. Concretely, it leads to 0.55 (full), 0.41 (rare) and 0.59 (non-rare) mAP improvements. In contrast, Contextualized Prompt Learning (CPL) achieves promising mAP increases in the full (0.94) and non-rare (1.23) settings. However, there is a slight decrease of 0.07 in mAP for the rare category. Fortunately, it can be effectively compensated by integrating the MRM module. When combined, they form a comprehensive model that delivers outstanding overall performance, notably achieving an impressive mAP boost of 1.50 in rare cases.
@@
Multivariate Relationship Modeling. The modeled binary connection is inherently included in the baseline, which cannot be removed. Therefore, we focus on unary association (UA) and ternary relationship (TR) modeling. As shown in the “MRM” column in Table II, both are effective in improving interaction recognition. For the Full, Rare and Non-rare metrics, associating the textual information in unary modeling can improve the mAP by 0.31, 0.44 and 0.26, respectively. Meanwhile, the sole use of ternary relationship modeling boost the mAP by 0.49, 0.34, and 0.53. The best performance is achieved when both of them are combined.
@@
Contextualized Prompt Learning. The proposed Contextual Decoder is crucial in CPL to refine textual features derived from learnable prompts. It comprises two blocks, namely the global block (GB) and the regional block (RB). As illustrated in the “CPL” column in Table II, both layers individually enhance HOI performance, and their combined use results in the most significant improvements.
@@
Moreover, those predefined parameters in CPL also need to be carefully analyzed. In Table III (a), we vary the length of learnable tokens [[ACT]] in our prompt template “A photo of a person [[ACT]] {object}”. The best performance is achieved with 4 tokens, while using more than this could potentially increase the complexity of learning. Instead of using the hand-crafted [[PREFIX]] tokens like “a photo of a”, we can use learnable ones, resulting in the format “[[PREFIX]] person [[ACT]] {object}”. As shown in Table III (b), switching to learnable [[PREFIX]] tokens appears to be less effective. This suggests that having too much flexibility in prompts might be suboptimal, aligning with the findings presented in Table III (a). Finally, as given in Table III (c), the performance is not particularly sensitive to changes in the CPL weight (i.e., $\beta$).
@@
Figure 5: Qualitative results on HICO-Det test set with fine-tuned DETR-R50 as the object detector. Bounding boxes of humans and objects are drawn with blue and green boxes. The textual annotation below the figure represents the ground truth. N/I denotes no interaction.
@@
Figure 6: The visualization of the cross-attention maps on a subset of images from Fig. 5. Bounding boxes of tools are drawn with red boxes. Best viewed in color.
IV-D Qualitative Results
As shown in Fig. 5, we present qualitative comparisons between our method and PViC [31] on the HICO-Det dataset. In the first row of the figure, we present three examples involving tool-related interactions. For actions related to “hit”, our model demonstrates more effective recognition of the “hit” interaction. Specifically, in the leftmost image of the first row, PViC tends to predict the interaction as “hold” or “carry”, whereas guided by tool-related semantics, our model correctly identifies the interaction as “fill”. It can be observed that our method performs robust in tool-related interaction cases and can consistently deliver more satisfying results. In the second row, we select examples of tool-irrelevant interactions for qualitative comparison. However, as illustrated by the rightmost example involving the action “paint”, painting typically requires a tool such as a brush. Since the predefined object detector cannot recognize brushes (as “paintbrush” is not among the 80 COCO categories), we did not include this sample in the HICO-Det-HTO dataset. In this example, our model assigns a higher score proportion to the “paint” interaction, while PViC incorrectly classifies the interaction mostly as “inspect”. In the two rightmost examples, PViC exhibits a tendency to classify the interactions as “hold”, while our approach successfully recognizes the more accurate interactions, namely “eat” and “hug”.
Moreover, as shown in Fig.6, we visualize attention maps corresponding to some of the images in Fig.5. Specifically, the first three examples illustrate the attention maps of the tool-related images in the first row of Fig.5, where the attention is taken from the last layer of the ternary decoder. We observe that, compared to PViC, our model places greater focus on the tools involved in the interactions. The last image in Fig.6 corresponds to the leftmost image in the second row of Fig.5. As previously mentioned, the brush in this image cannot be detected by the object detector, and thus cannot be recognized as a tool. Consequently, the ternary token cannot be constructed, and no attention map is available from the ternary decoder. Instead, we visualize the attention map from the last layer of the pairwise decoder, and it shows that our model still attends to the brush held in the person’s hand.
During our experiments, we observed that existing datasets such as HICO-Det are not well-suited for annotating certain tool-related categories. For example, in some images depicting “a person hitting a baseball”, but the annotations are often limited to actions such as “person swinging a bat” or “person holding a bat”, which fail to capture the full semantics of the interaction. Moreover, the number of images involving measurable tool-related interactions is relatively small, making it difficult to effectively leverage tool-related information. In future work, we plan to expand the dataset by collecting more images, either from real-world sources or by generating synthetic images using generative models such as diffusion models.
TABLE IV: Zero-shot comparisons with SOTA methods on HICO-Det.
Method Type Unseen Seen Full
CMMP [14] RF-UC 29.45 32.87 32.18
EZ-HOI [34] RF-UC 29.02 34.15 33.13
LAIN [43] RF-UC 31.83 35.06 34.41
HOLa [44] RF-UC 30.61 35.08 34.19
Ours RF-UC 31.73 36.61 35.63
CMMP [14] NF-UC 32.09 29.71 30.18
EZ-HOI [34] NF-UC 33.66 30.55 31.17
LAIN [43] NF-UC 36.41 32.44 33.32
HOLa [44] NF-UC 35.25 31.64 32.36
Ours NF-UC 28.95 32.59 31.87
CMMP [14] UO 33.76 31.15 31.59
EZ-HOI [34] UO 33.28 32.06 32.27
LAIN [43] UO 37.88 33.55 34.27
HOLa [44] UO 36.45 33.02 33.59
Ours UO 36.74 34.53 34.90
CMMP [14] UV 26.23 32.75 31.84
EZ-HOI [34] UV 25.10 33.49 32.32
LAIN [43] UV 28.96 33.80 33.12
HOLa [44] UV 27.91 35.09 34.09
Ours UV 13.05 36.05 32.83
IV-E Discussion
@@
Model Complexity. Considering that the introduction of the ternary decoder and contextual decoder introduces additional trainable parameters, we conduct a comprehensive comparison with several recent one-stage and two-stage HOI models. Specifically, we compare our method with EZ-HOI [15], CMMP [14], and ADA-CM [35], all of which adopt ViT-L/14 as the backbone. We also include comparisons with our baseline PViC [31] and the one-stage method HOICLIP [34].
@@
Zero-shot Setting. Table IV compares the zero-shot HOI detection performance of our method with four SOTA approaches (CMMP, EZ-HOI, LAIN, HOLa) on HICO-Det across four scenarios (RF-UC, NF-UC, UO, UV), evaluating Unseen, Seen, and Full metrics. Since the experiment is conducted under a zero-shot setting, we freeze the last three classifiers and adopt the Verb Adapter from HOICLIP[34] as the weights for these three classifiers. Our method exhibits notable strengths: it ranks first in the Seen metric across all scenarios (e.g., 36.61% in RF-UC, 36.05% in UV) and secures top positions in the Full metric for RF-UC (35.63%) and UO (34.90%), with competitive Unseen performance in these two scenarios (31.73% for RF-UC, 36.74% for UO). However, it has room for improvement: in NF-UC, its Unseen score (28.95%) drags the Full metric to 31.87. In the unseen verb (UV) task, the low Unseen score (13.05%) restricts the Full metric to only 32.83%. The poor performance of unseen verbs stems from two key issues: tool-related ones like “stab” and “swing” are often misclassified as familiar actions such as “hit” or “serve” due to imprecise feature links, while abstract verbs like “inspect” and “install” lack consistent connections to known actions. Their vague boundaries and context-dependent use further hinder knowledge transfer. To fix this, future work should refine tool-action representations to capture subtle differences and integrate goal information for abstract verbs.
V Conclusion
In this paper, we have proposed CRL, a two-stage HOI detection framework that addresses two key limitations of existing methods: the lack of tool-mediated modeling and the inability to integrate visual features into prompt learning, limiting their capacity to capture context-dependent relationships. Our approach has shown notable effectiveness in challenging HOI scenarios with complex tool-assisted interactions and ambiguous contextual cues, highlighting the benefits of multivariate relationship modeling and contextualized prompt learning. In future work, we plan to explore fine-grained affordance-guided modeling and broaden generalization to open-world HOI scenarios.